An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 808 hours and a standard deviation of 93 hours. Find the probability that a random samples of 19 bulbs will have an average life of greater than 799 hours.


1
Expert's answer
2022-02-10T17:51:20-0500

Let "X=" an average life: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=808\ h, \sigma=93\ h, n=19."


"P(X>799)=1-P(X\le 799)"

"=1-P(Z\le \dfrac{799-808}{93/\sqrt{19}})"

"\approx 1-P(Z\le-0.421829)"

"\approx0.663425"